An equilateral triangle has a side length of 4 ft. What is the area of the triangle?
Answer by Boreal(15235)

Draw this and the altitude is a perpendicular bisector of the base.
that right triangle has base 2, alt h and an angle of 60 degrees
sin 60=alt/4
sqrt(3)/2=alt/4
altitude is 2 sqrt(3) feet
area is 1/2 bh=1/2*2sqrt(3)*4=4 sqrt(3) feet
the general formula is s^2/4* sqrt(3)